Can you help me find the x- and y-intercepts and graph the equation plotting the intercepts? -2x + 3y = -15
Thank you so much!

Answer by mananth(16949)

-2x + 3y = -15
3y=2x-15
y =2x/3-5
y intercept, x=0
when x=0 y = -5 (0,-5)
x Intercept
when y=0, x =15/2 ( 15/2,0)
Plot the points draw the line